OpenStudy (dancergirl45):

Which of the following options is an equivalent function to f(x) = 2(5)^2x? f(x)=50^x f(x)=100^x f(x)=2(25)^x f(x)=4(25)^x

OpenStudy (anonymous):

so d is your answer

OpenStudy (dancergirl45):

That is what i thought

OpenStudy (dancergirl45):

so it is right

OpenStudy (dancergirl45):

@jayshon54rogers

OpenStudy (trojanpoem):

f(x) = 2 * (5)^2x = 2 * ((5)^2)^x = 2 (25)^x

OpenStudy (dancergirl45):

so the answer is D

OpenStudy (trojanpoem):

1) f(x)=50^x 2) f(x)=100^x 3) f(x)=2(25)^x <<< 4) f(x)=4(25)^x

OpenStudy (dancergirl45):

ok C thank you!!

OpenStudy (trojanpoem):

Any time.

OpenStudy (dancergirl45):

can you help me with another i will open a new question

OpenStudy (trojanpoem):

Sure.